Hi, I'm looking for help with solving this problem 8(x-4)+10=18

Answer by JulietG(1812)

Review Order of Operations for additional problems.
The first equation is to distribute the 8 throughout the parentheses.
8(x-4)+10=18
8x-32+10=18
Now add
8x-22=18
Add 22 to each side (you're trying to isolate the x)
8x = 40
Divide each side by 8
x = 5